Could someone please show me how to solve by substitution method?
8x - 2y = -70
3x + 26 = y

Found 2 solutions by ewatrrr, rwm:
Answer by ewatrrr(24785)


Hi
y = 3x+26
8x - 2y = -70
8x-2(3x+26) = -70 |substituting for y
8x - 6x - 52 = -70
2x = -18
x = -9 and y = -1 (-27+26)
CHECKING our Answer***
8x - 2y = -70
-72 + 2 = -70

Answer by rwm(914)

plug in 3x+26 for y in the first equation
8x-2*(3x+26)=-70
now solve for x
8x-6x-52=-70
2x=-18
x=-9
now plug 9 in for x in either equation
3*(-9)+26=y
-27+26=y
-1=y